# Impact of Small-Scale Gravity Waves on Tracer Transport

### v1

@@ 4.2 Parameterization of the Leading-order Gravity Wave Tracer Flux Convergence
Figure 1: Gravity wave-induced tracer flux convergence $\mathcal{Q}^{(0)}$ (top row) and change in the large-scale tracer distribution $\Delta\left\langle{\psi}\right\rangle$ (bottom row) from a three-dimensional gravity wave packet, at $t=1500\text{min}$; (left column) 3DWR: wave-resolving reference simulation; (center column) 3DCR: coarse-resolution simulation with parameterized gravity wave and including $\mathcal{Q}^{(0)}$; (right column) 3DCRno: coarse-resolution simulation with $\mathcal{Q}^{(0)}=0$. The solid and dashed contour lines illustrate the regions with positive and negative values, respectively.
@@ 4.3 Parameterization of the Next-order Gravity Wave Tracer Flux Convergence
From the extension of MS-GWaM to calculate $\mathcal{Q}^{(1)}$, we assume a locally monochromatic wave packet. Dropping the index $\alpha$ from the equations summarized in Section 3, the wave energy is given
@@
Within MS-GWaM, the wavenumber integral is calculated by summing the contributions from all ray volumes within a PincFlow finite volume cell. As discussed in Section 2, we can determine both the leading- and next-order wave amplitudes from the buoyancy amplitude $\hat{b}=\left|\hat{b}\right|e^{i\varLambda}$, allowing us to parameterize the next-order tracer flux convergence. Numerically, all required temporal and spatial derivatives are approximated using simple finite differences. It is important to note that in the present monochromatic case, Equation 68 predicts that the large-scale phase remains invariant.
@@
where Equation 90 is used in coarse-resolution simulations with parameterized gravity waves through MS-GWaM, while wave-resolving simulations solve Equation 70.
We conduct a wave-resolving simulation (2DWR) as a reference simulation and a simulation with parameterized gravity waves (2DCR). To emphasize the impact of $\mathcal{Q}^{(1)}$, we also conduct an additional coarse-resolution simulation with parameterized waves where we set $\mathcal{Q}^{(1)}=0$ (2DCRno). In the case of 2DWR, the resolution is 16 grid points per wavelength in the $x$-direction and 10 grid points per vertical wavelength in the $z$-direction. For 2DCR and 2DCRno, we choose one grid point per wavelength in both the horizontal and vertical directions.
@@
from which all remaining perturbations in the atmospheric variables are calculated using the polarization relations. Since the wave packet is meant to propagate upwards, we select the positive branch of the frequency with $k_{0}>0$ and $m_{0}<0$ for the horizontal and vertical wavenumbers, respectively. As we are interested only in the second-order gravity wave effects, and not in turbulent mixing induced by breaking gravity waves, we set $a_{0}=0.1$ so that the gravity wave packet remains far from static instability throughout the simulation. The initial large-scale tracer distribution is linearly increasing with altitude, as given in Equation 84.
We calculate the large-scale variables from the 2DWR results by averaging over one wavelength in the horizontal direction. From there, we can calculate the perturbations in the atmospheric variables, as given in Equations 85 and 86. Using these, we can then calculate the tracer flux convergence
@@
The results of 2DWR, 2DCR, and 2DCRno over a simulation time of 60 min are visualized in Figure 2. Both the next-order tracer flux convergence $\mathcal{Q}^{(1)}$ and the change in the large-scale tracer distribution calculated in 2DWR are replicated by the parameterized gravity wave simulation. We observe that the gravity wave causes a vertical tracer flux towards the center of the wave packet. In the case of 2DCRno, where $\mathcal{Q}^{(1)}=0$, we observe no change in the large-scale tracer distribution, as there is no large-scale vertical wind (not shown). We can therefore conclude that the $\Delta\left\langle{\psi}\right\rangle$ seen in the results for 2DWR and 2DCR are accurately parameterized in 2DCR.
Figure 2: Time evolution of the gravity wave-induced tracer flux convergence $\mathcal{Q}^{(1)}$ (top row) and change in the large-scale tracer distribution $\Delta\left\langle{\psi}\right\rangle$ (bottom row) from a two-dimensional gravity wave packet; (left column) 2DWR: wave-resolving reference simulation; (center column) 2DCR: coarse-resolution simulation with parameterized gravity wave and including $\mathcal{Q}^{(1)}$; (right column) 2DCRno: coarse-resolution simulation with $\mathcal{Q}^{(1)}=0$. The solid and dashed contour lines illustrate the regions with positive and negative values, respectively.
@@ 5 Summary and conclusions
Our findings reveal the leading-order and next-order contributions of small-scale gravity waves to large-scale tracer transport. The leading-order impact, induced by inertial gravity waves, acts perpendicular to both the large-scale tracer gradient and the wave number. The next-order impact becomes particularly significant at lower latitudes, where the Coriolis parameter becomes small and therefore also the leading-order impact, due to its non-zero contribution in the absence of rotation. These impacts scale as $\mathcal{O}\left(\varepsilon^{2\gamma+1}\right)$ and $\mathcal{O}\left(\varepsilon^{2\gamma+2}\right)$, respectively, when compared to the large-scale wind advection ($\mathcal{O}(1)$), with $\gamma=0,1,2$ corresponding to large-amplitude, weakly nonlinear, and quasi-linear waves. Additionally, we demonstrated that gravity wave effects can be calculated from the leading-order buoyancy amplitude and gravity wave energy. We extended a gravity wave parameterization model to calculate these effects. Validation of these results was performed using two idealized wave packet test cases, where comparisons with wave-resolving reference simulations confirmed the accuracy of the approach.

### v2

@@ 4.2.2 Wave packet in an atmosphere with a jet
Figure 2: Gravity wave-induced tracer flux convergence $\mathcal{Q}^{(0)}$ (top row) and change in the large-scale tracer distribution $\Delta\left\langle{\psi}\right\rangle$ (bottom row) from a three-dimensional gravity wave packet with a jet centered at $z=$10\text{\,}\mathrm{km}$\$, at $t=$6\text{\,}\mathrm{h}$\$; (left column) 3DWR: wave-resolving reference simulation; (center column) 3DCR: coarse-resolution simulation with parameterized gravity wave and including $\mathcal{Q}^{(0)}$; (right column) 3DCRno: coarse-resolution simulation with $\mathcal{Q}^{(0)}=\mathcal{Q}^{(1)}=0$. The solid and dashed contour lines illustrate the regions with positive and negative values, respectively.
@@ 4.3 Parameterization of the Next-order Gravity Wave Tracer Flux Convergence
Figure 3: Time evolution of the gravity wave-induced tracer flux convergence $\mathcal{Q}^{(1)}$ (top row) and change in the large-scale tracer distribution $\Delta\left\langle{\psi}\right\rangle$ (bottom row) from a two-dimensional gravity wave packet; (left column) 2DWR: wave-resolving reference simulation; (center column) 2DCR: coarse-resolution simulation with parameterized gravity wave and including $\mathcal{Q}^{(1)}$; (right column) 2DCRno: coarse-resolution simulation with $\mathcal{Q}^{(0)}=\mathcal{Q}^{(1)}=0$. The solid and dashed contour lines illustrate the regions with positive and negative values, respectively.
@@
From the extension of MS-GWaM to calculate $\mathcal{Q}^{(1)}$, we assume a locally monochromatic wave packet. Dropping the index $\alpha$ from the equations summarized in Section˜3, the wave energy is given
@@
Within MS-GWaM, the wavenumber integral is calculated by summing the contributions from all ray volumes within a PincFlow finite volume cell. As discussed in Section˜2, we can determine both the leading- and next-order wave amplitudes from the buoyancy amplitude $\hat{b}=\left|\hat{b}\right|e^{i\varLambda}$, allowing us to parameterize the next-order tracer flux convergence. Numerically, all required temporal and spatial derivatives are approximated using simple finite differences. It is important to note that in the present monochromatic case, Equation˜69 predicts that the large-scale phase remains invariant.
@@
where Equation˜94 is used in coarse-resolution simulations with parameterized gravity waves through MS-GWaM, while wave-resolving simulations solve Equation˜73.
We conduct a wave-resolving simulation (2DWR) as a reference simulation and a simulation with parameterized gravity waves (2DCR). To emphasize the impact of $\mathcal{Q}^{(1)}$, we also conduct an additional coarse-resolution simulation with parameterized waves where we set $\mathcal{Q}^{(1)}=0$ (2DCRno). In the case of 2DWR, the resolution is 16 grid points per wavelength in the $x$-direction and 10 grid points per vertical wavelength in the $z$-direction. For 2DCR and 2DCRno, we choose one grid point per wavelength in both the horizontal and vertical directions. Furthermore, as the atmosphere is nonrotating, $\mathcal{Q}^{(0)}=0$ for both 2DCR and 2DCRno.
@@
from which all remaining perturbations in the atmospheric variables are calculated using the polarization relations. Since the wave packet is meant to propagate upwards, we select the positive branch of the frequency with $k_{0}>0$ and $m_{0}<0$ for the horizontal and vertical wavenumbers, respectively. As we are interested only in the second-order gravity wave effects, and not in turbulent mixing induced by breaking gravity waves, we set $a_{0}=0.1$ so that the gravity wave packet remains far from static instability throughout the simulation. The initial large-scale tracer distribution is linearly increasing with altitude, as given in Equation˜87.
We calculate the large-scale variables from the 2DWR results by averaging over one wavelength in the horizontal direction. From there, we can calculate the perturbations in the atmospheric variables, as given in Equations˜88 and 89. Using these, we can then calculate the tracer flux convergence
@@
The results of 2DWR, 2DCR, and 2DCRno over a simulation time of 60 min are visualized in Figure˜3. Both the next-order tracer flux convergence $\mathcal{Q}^{(1)}$ and the change in the large-scale tracer distribution calculated in 2DWR are replicated by the parameterized gravity wave simulation. We observe that the gravity wave causes a vertical tracer flux towards the center of the wave packet. In the case of 2DCRno, where $\mathcal{Q}^{(1)}=0$, we observe no change in the large-scale tracer distribution, as there is no large-scale vertical wind (not shown). We can therefore conclude that the $\Delta\left\langle{\psi}\right\rangle$ seen in the results for 2DWR and 2DCR are accurately parameterized in 2DCR.
@@ 5 Summary and conclusions
Our findings reveal the leading-order and next-order contributions of small-scale gravity waves to large-scale tracer transport. The leading-order impact, induced by inertial gravity waves, acts perpendicular to both the large-scale tracer gradient and the wave number. The next-order impact becomes particularly significant at lower latitudes, where the Coriolis parameter becomes small and therefore also the leading-order impact, due to its non-zero contribution in the absence of rotation. These impacts scale as $\mathcal{O}\left(\varepsilon\right)$ and $\mathcal{O}\left(\varepsilon^{2}\right)$, respectively, when compared to the large-scale wind advection ($\mathcal{O}(1)$). Additionally, we demonstrated that gravity wave effects can be calculated from the leading-order buoyancy amplitude and gravity wave energy. We extended a gravity wave parameterization model to calculate these effects. Validation of these results was performed using three idealized wave packet test cases, where comparisons with wave-resolving reference simulations confirmed the accuracy of the approach.
@@ acknowledgements
Appendix A Generalized Derivation of Gravity Wave-Induced Tracer Fluxes
In Sections˜2 and 3, we derived expressions for the gravity wave-induced tracer fluxes (Equations˜66 and 67), summarized in dimensional form, under the simplifying assumption of a locally monochromatic, large-amplitude gravity wave in a weakly stratified atmosphere ($\delta=1$). In this appendix, we extend the derivation to a more general setting. Specifically, we consider strongly stratified atmosphere ($\delta=0$), weakly nonlinear and quasi-linear gravity wave fields, and spectra of waves including higher harmonics.
For a spectrum of gravity waves with $A$ wavenumbers, the atmospheric variables can be decomposed into three contributions: the reference atmosphere, the large-scale flow, and the wave field:
The wave terms are proportional to $e^{i\beta\phi_{\alpha}/\varepsilon}$. Here, $\alpha$ indexes distinct waves, each with its own phase function $\phi_{\alpha}$. The harmonic index $\beta$ refers to integer multiples of the same phase, so that $\beta$-th harmonic of wave $\alpha$ has phase $\beta\phi_{\alpha}$. In other words, different phase structures belong to different waves (different $\alpha$), while harmonics of a given wave share the same underlying phase function. The corresponding wavenumbers and frequencies are denoted by $\beta\mathbf{k}_{\alpha}=\beta\nabla_{\mathbf{X}}\phi_{\alpha}$ and $\beta\omega_{\alpha}=-\beta\partial_{\mathrm{T}}\phi_{\alpha}$, respectively.
The scaling factor $\varepsilon^{\gamma}$ distinguishes three regimes:
  • Large-amplitude waves ($\gamma=0$): Locally monochromatic case with $A=1$.
  • Weakly nonlinear waves ($\gamma=1$): Spectrum of waves allows ($A\geq 1$), but wave-wave interactions affecting the gravity waves themselves are neglected.
  • Quasi-linear waves ($\gamma=2$): No restrictions.
We impose no amplitude assumptions on the tracer field, whose magnitude is determined directly from the governing equations.
These assumptions are essential to keep the governing equations solvable. Higher-order amplitude terms introduce wave-wave interactions, which render the system nonlinear in the wave amplitudes and prohibitively complex. Since our goal is to construct a system of equations suitable for parameterization, we exclude such terms. This simplification is also physically justified, as the essential impact of gravity waves on tracer fluxes is preserved.
A summary of the indices $\alpha$, $\beta$, $\gamma$, and $\delta$ is given in Table˜3.
The derivations of the tracer flux convergences $\mathcal{Q}^{(0)}$ and $\mathcal{Q}^{(1)}$ proceed analogously to Section˜2. Here we present only the generalized form and highlight the main differences. The leading-order tracer convergence becomes:
The next-order flux convergence is:
Only the fundamental wave ($\beta=1$) contributes to the flux convergence, since $\hat{\mathbf{V}}_{\alpha\beta}^{(0)}=0$ for all $\beta>1$ [2]. Consequently, the tracer amplitude $\hat{\Psi}_{\alpha\beta}^{(\gamma+1)}$ vanishes for higher harmonics as
$i\beta\hat{\omega}_{\alpha}\hat{\Psi}_{\alpha\beta}^{(\gamma+1)}=\hat{\mathbf{V}}_{\alpha\beta}^{(0)}\cdot\nabla_{\mathbf{X}}\left\langle{\underline{\Psi}}\right\rangle\;.$ (104)
The intrinsic frequency is denoted by $\hat{\omega}_{\alpha}=\omega_{\alpha}-\mathbf{k}_{h,\alpha}\cdot\left\langle{\underline{\mathbf{V}}}\right\rangle$. To obtain $\mathcal{Q}^{(0)}$, we employ the wind amplitude relations, modified by replacing in Equations˜21 and 22 the amplitude $\hat{\mathbf{V}}^{(m)}$ with $\hat{\mathbf{V}}_{\alpha 1}^{(m)}$ for $m=0,1$, and $\mathbf{k}$ with $\mathbf{k}_{\alpha}$. For $\mathcal{Q}^{(1)}$, we use the next-order tracer amplitude relation
As noted, wave-wave interactions in the weakly nonlinear regime ($\gamma=1$) would appear as triad terms in Equation˜105. These are neglected here to keep the problem tractable. Similarly, in the strongly stratified case ($\delta=0$), the correction terms
$i\frac{c_{V}}{R}\frac{i\hat{\omega}_{\alpha}}{\underline{\overline{\Pi}}}\hat{\Pi}^{(0)}_{\alpha 1}-\frac{\left\langle{\underline{\Pi}}\right\rangle}{\overline{\underline{\Pi}}}\frac{1}{\left\langle{P}\right\rangle}\nabla\cdot\left(\left\langle{P}\right\rangle\hat{\mathbf{V}}_{\alpha 1}^{(0)}\right)+\frac{\left\langle{\underline{\Pi}}\right\rangle}{\overline{\underline{\Pi}}}\frac{1}{\overline{P}}\nabla_{\mathbf{X}}\cdot\left(\overline{P}\hat{\mathbf{V}}_{\alpha 1}^{(0)}\right)\;,$ (106)
appear on the right-hand side of the relation for the wind amplitude
$i\mathbf{k}_{\alpha}\cdot\hat{\mathbf{V}}_{\alpha 1}^{(2)}=-\frac{1}{\overline{P}}\nabla_{\mathbf{X}}\cdot\left(\overline{P}\hat{\mathbf{V}}_{\alpha 1}^{(1)}\right)\;,$ (107)
with $\left\langle{P}\right\rangle=\left\langle{\rho}\right\rangle\left\langle{\theta}\right\rangle$, which we neglect for tractability. For a spectrum of waves, the leading-order flux convergence can be written as
$\mathcal{Q}^{(0)}=-\frac{\varepsilon^{(2\gamma+1)}}{\overline{\rho}}\sum_{\alpha=1}^{A}\nabla_{\mathbf{X}}\cdot\left(\frac{f_{0}}{\hat{\omega}_{\alpha}}\frac{m_{\alpha}}{\varepsilon^{5-\delta}k_{h,\alpha}^{2}+m_{\alpha}^{2}}\mathcal{A}_{\alpha}\mathbf{k}_{\alpha}\times\nabla_{\mathbf{X}}\left\langle{\underline{\Psi}}\right\rangle\right)\;,$ (108)
with wave action density $\mathcal{A}_{\alpha}=E_{w,\alpha}/\hat{\omega}_{\alpha}$, wave energy $E_{w,\alpha}$
$E_{w,\alpha}=\overline{\rho}\frac{\left|\hat{B}_{\alpha 1}^{(0)}\right|^{2}}{2N_{0}^{2}}\frac{N_{0}^{2}k_{h,\alpha}^{2}+f_{0}^{2}m_{\alpha}^{2}}{N_{0}^{2}k_{h,\alpha}^{2}}\;,$ (109)
and the leading-order buoyancy amplitude $\hat{B}_{\alpha 1}^{(0)}=\hat{\Theta}_{\alpha 1}^{(0)}/\overline{\underline{\Theta}}$. For the calculation of $\mathcal{Q}^{(1)}$, we first decompose the buoyancy amplitude
$\hat{B}_{\alpha 1}^{(0)}=\left|\hat{B}_{\alpha 1}^{(0)}\right|e^{i\varLambda_{\alpha}}\;.$ (110)
The absolute magnitude follows from the wave energy in Equation˜109, while the large-scale phase $\varLambda_{\alpha}$ evolves according to the prognostic equation:
$\begin{split}\left(\partial_{\mathrm{T}}+\mathbf{c}_{g,\alpha}\cdot\nabla_{\mathbf{X}}\right)\varLambda_{\alpha}=&-\frac{f_{0}m_{\alpha}}{2\left|\mathbf{k}_{\alpha}\right|^{2}\hat{\omega}_{\alpha}}\left(\mathbf{k}_{\alpha}\times\nabla_{\mathbf{X}}\right)\left\langle{\underline{\mathbf{U}}}\right\rangle\\ &+\frac{|\mathbf{k}_{\alpha}|^{2}}{2m_{\alpha}\left(N_{0}^{2}-\varepsilon^{5-\delta}f_{0}^{2}\right)}\left[\frac{1}{\overline{\underline{\Theta}}}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Theta}}\right\rangle-\varepsilon^{\delta}N_{0}^{2}\frac{c_{p}}{R}\overline{\underline{\Theta}}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Pi}}\right\rangle\right]\\ &-\frac{f_{0}\hat{\omega}_{\alpha}m_{\alpha}^{2}}{k_{h,\alpha}^{2}\left(N_{0}^{2}-\varepsilon^{5-\delta}f_{0}^{2}\right)}\left[\left(l_{\alpha}\partial_{\mathrm{X}}-k_{\alpha}\partial_{\mathrm{Y}}\right)\hat{\omega}_{\alpha}+l_{\alpha}\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)k_{\alpha}-k_{\alpha}\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)l_{\alpha}\right]\;,\end{split}$ (111)
with group velocity $\mathbf{c}_{g,\alpha}=\nabla_{\mathbf{k}}\omega_{\alpha}$ and intrinsic group velocity $\hat{\mathbf{c}}_{g,\alpha}=\nabla_{\mathbf{k}}\hat{\omega}_{\alpha}$. Next, we solve the matrix equation:
$M_{\alpha 1}\mathbf{Z}_{\alpha 1}^{(1)}=\mathbf{R}_{\alpha 1}^{(1)}\;,$ (112)
with matrix
next-order wave amplitude vector
and right-hand side vector
The components of $\mathbf{R}_{\alpha 1}^{(1)}$ are
where $\mathbf{R}_{\mathbf{u},\alpha 1}^{(1)}=\left(R_{u,\alpha 1}^{(1)}\>,\>R_{v,\alpha 1}^{(1)}\right)^{\mathrm{T}}$. The leading-order wave amplitudes are obtained from $\hat{B}_{\alpha 1}^{(0)}$ using the polarization relations
Finally, with Equation˜105, the next-order flux convergence in Equation˜103 can be evaluated.
In dimensional form, the leading-order tracer flux convergence is:
while the next-order flux convergence is:
$\begin{split}\mathcal{Q}^{(1)}=-\frac{1}{2\overline{\rho}\overline{\theta}}\sum_{\alpha=1}^{A}&\nabla\cdot\mathfrak{R}\left[\overline{\rho}\overline{\theta}\left(\mathbf{v}_{\alpha}^{\prime}\psi_{\alpha}^{\prime\prime*}+\mathbf{v}_{\alpha}^{\prime\prime}\psi_{\alpha}^{\prime*}\right)\right]\;.\end{split}$ (124)
The calculation of $\mathcal{Q}^{(1)}$ proceeds sequentially:
  1. 1.
    Solve for the leading-order buoyancy amplitude $b_{\alpha 1}^{\prime}=\left|b_{\alpha 1}^{\prime}\right|e^{i\varLambda_{\alpha}}$ from the wave energy equation
    and integrate the prognostic equation for $\varLambda_{\alpha}$ to obtain the large-scale phase:
    $\begin{split}\left(\partial_{t}+\mathbf{c}_{g,\alpha}\cdot\nabla\right)\varLambda_{\alpha}=&-\frac{fm_{\alpha}}{2\left|\mathbf{k}_{\alpha}\right|^{2}\hat{\omega}_{\alpha}}\left(\mathbf{k}_{\alpha}\times\nabla\right)\left\langle{\mathbf{u}}\right\rangle\\ &+\frac{|\mathbf{k}_{\alpha}|^{2}}{2m_{\alpha}\left(N^{2}-f^{2}\right)}\left[\frac{g}{\overline{\theta}}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla\right)\left\langle{\theta}\right\rangle-\frac{N^{2}}{g}c_{p}\overline{\theta}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla\right)\left\langle{\pi}\right\rangle\right]\\ &-\frac{f\hat{\omega}_{\alpha}m_{\alpha}^{2}}{k_{h,\alpha}^{2}\left(N^{2}k_{h,\alpha}^{2}+f^{2}m_{\alpha}^{2}\right)}\left[\left(l_{\alpha}\partial_{x}-k_{\alpha}\partial_{y}\right)\hat{\omega}_{\alpha}+l_{\alpha}\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)k_{\alpha}-k_{\alpha}\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)l_{\alpha}\right]\;.\end{split}$ (126)
  2. 2.
    Compute the leading-order wind and Exner pressure amplitudes using the polarization relations:
  3. 3.
    Evaluate the right-hand side terms:
    and solve the matrix equation for $\mathbf{Z}^{\prime\prime}_{\alpha}$:
  4. 4.
    Determine the leading- and next-oder tracer amplitudes from:
    and
  5. 5.
    Substitute these amplitudes and the corresponding wind fields into Equation˜124 to obtain $\mathcal{Q}^{(1)}$.
This generalized derivation extends the results of Section˜2 and Section˜3 to cover spectra of gravity waves, large-amplitude, weakly nonlinear, and quasi-linear regimes, and both weakly and strongly stratified atmospheres. While simplified for clarity, the formulation provides a flexible basis for parameterizing tracer fluxes under realistic atmospheric conditions.
Table 3: Summary of the indices that represent various wave cases and atmospheric conditions.
[…]
$\alpha$ Wave modes: indicates the wave in a spectrum of waves with $A$ wavenumbers $\{1...A\}$
$\beta$ Higher harmonics of the wave
11: basic wave
>1>1: higher harmonic
$\gamma$ Measure of non-linearity of the gravity waves
0: high-amplitude wave
11: weakly nonlinear
22: quasi-linear
$\delta$ Stratification of the reference atmosphere
0: strong
11: weak